OpenStudy (anonymous):

Find the equation x-intercept= -1/5 slope=-3

OpenStudy (rane):

first u have to know what kind of equation it is?

OpenStudy (anonymous):

this is a linear equation... from the given values of x intercepts and the slope: note that () from , slope=change in y/change in x m=y2 - y1/x2-x1 -3=y2-0/x2-(-1/5) -3=y2/x2+1/5 -3(x2+1/5)=y2 -3x2-3/5=y2 note: just assume x2 and y2 to be x and y valuables -3x -3/5=y -3x-y=3/5 3x+y=-3/5 this is the linear equation..